### v1

# A voice for minorities: diversity in approval-based committee elections under incomplete or inaccurate information

Abstract
We study diversity in approval-based committee elections with incomplete or inaccurate information. As standard in the literature on approval-based multi-winner voting, we define diversity according to the maximum coverage problem, which is known to be NP-complete, with a best attainable polynomial time approximation ratio of $1-1/\operatorname{\text{e}}$. In the incomplete information model, voters can vote on only a small portion of the candidates. We suggest a greedy algorithm and a local search algorithm that query voters and use the query responses to approximate the total population’s opinion. For both algorithms, we prove an upper bound on the number of queries required to get a close to $(1-1/\operatorname{\text{e}})$-approximate solution with high probability. We also provide a lower bound for the query complexity of non-adaptive algorithms, that cannot adapt their querying strategy to readily obtained information. In the inaccurate information setting, voters’ responses are corrupted with a probability $p\in(0,\frac{1}{2})$. We provide both an upper and a lower bound for the number of queries required to attain a $(1-1/\operatorname{\text{e}})$-approximate solution with high probability. Finally, using real data from Polis, we see that our algorithms perform remarkably better than the theoretical results suggest, both with incomplete and inaccurate information.
1 Introduction
Consensus has grown among political scientists, democracy practitioners, and decision makers alike, that effective involvement of citizens in policy decisions should be regarded as a high priority for democratic institutions at all levels, from local, to national, to regional [18, 24]. [^1] See also initiatives such as “Conference on the Future of Europe” (2021-22) and “European Democracy Shield” (2024). Many well-established democratic innovations [34], such as citizens’ assemblies and participatory budgeting, contribute to this civic participation goal. Yet, it is on digital democracy platforms in particular where research and development efforts are currently concentrating [6, 13], with the aim of providing citizens with effective tools that could support forms of democratic deliberation at scale [28, 25].
A wealth of this type of digital democracy tools have been developed and deployed around the world in the last decade: LiquidFeedback [3], Consul (recipient of the United Nations Public Service Award), YourPriorities, Decidim, Polis [30], to mention a few. [^2] See www.liquidfeedback.org; www.consulproject.org; www.citizens.is; www.decidim.org; www.pol.is. In different forms, and to a smaller or larger extent, all these applications provide users with the freedom to contribute statements, ideas, arguments to public deliberations on issues relevant to their communities and enable them to express their own opinions on the statements/ideas/arguments expressed by others. Supporting such processes online in large groups necessarily involves the implementation of algorithmic solutions for information-processing problems. One such problem concerns how to feed back to users information that can effectively summarize the current state of the deliberation: how does ‘the group’ think about the issues being deliberated upon?
One approach, which is followed for instance in Polis, to aggregate the current state of the deliberation into a comprehensive summarization of it, is by selecting sets of statements proposed by users, based on how much support the statements have elicited from other users. As [15] already notice, this approach can be conceptualized as an approval-based multi-winner voting problem [11, 21], as studied in computational social choice [2]: users ‘vote’ for other users’ statements by approving or disapproving them. Unlike in standard multi-winner approval voting, however, information about who supports which statements is sparse in online deliberation summarization, as no user can possibly express whether they agree or disagree with every single statement contributed by their peers. For the same reason, voters can become inaccurate in representing their true beliefs, when having to answer many questions. So, such a deliberation summarization problem can be thought of as an approval-based multi-winner voting problem in which ballots are incomplete or inaccurate. From this perspective, [15] focused on the problem of designing selection algorithms that could yield summarizations that meet specific forms of proportional representation proposed in the computational social choice literature [1], while querying users’ opinions as efficiently as possible.
Our work builds on the approach put forth in [15], but leaves proportional representation aside to focus on the property of diversity. Diversity has received much less attention than proportional representation in the multi-winner voting literature [21]. Yet, we argue diversity is especially important for deliberation summarization. Firstly, diverse summarizations would guarantee higher inclusivity, an attribute one could desire in and of itself. Secondly, a diverse summary gives a broad, informative view of ‘the people’s’ opinion. This can be valuable especially in deliberation summarization contexts, where the number of candidates may be very large. Thirdly, cognitive diversity is argued, e.g., by [18], to also be of epistemic advantage from a crowd-wisdom perspective. This is relevant because the selected statements typically serve as a basis for further discussions. We start from the established formalization of diversity in multi-winner voting based on the Chamberlin-Courant [7] score and the Maximum Coverage problem [26, 32]. Our focus thus lies on studying the computational problem of constructing summarizations that maximize the Chamberlin-Courant score, while querying users’ opinions as efficiently as possible.
2 Related work and our contribution
Diversity in multi-winner voting and the Maximum Coverage problem
Diversity in multi-winner voting is commonly defined following the characterization in the seminal work of Chamberlin and Courant in 1983 on representative deliberations and representative decisions [7]. Their most general definition uses voters’ ordinal preferences and, for any given committee, a voter’s representative in the committee is the member who ranks highest in the voter’s preference order. The committee score is obtained by summing over the individual voters’ scores, attributed to them according to the position of their representative in their preference order. In the case of approval ballots, maximizing the Chamberlin-Courant score simplifies to selecting a set of candidates that guarantees for as many voters as possible at least one candidate they approve of. In 2008, [29] have proven, using a reduction from Exact 3-Cover, the Chamberlin-Courant decision problem is NP complete, even for the version with approval ballots. Denoting candidates by the set of their approving voters, the approval-based problem is also equivalent to selecting a set of candidates that maximizes the coverage of the set of voters. This one-to-one correspondence between Chamberlin-Courant score-based diversity in approval-based multi-winner voting, and the Maximum Coverage problem, was first studied in [32].
An instance of the Maximum Coverage problem, Max Cover in short, consists of a set $V$ of $n$ elements, a weight function $w:V\to\mathbb{R}_{>0}$, a family $C\subseteq\mathcal{P}(V)$ of $m$ subsets, and a number $k$. The goal then is to select $k$ sets $c_{i}\in C$ so that $w(\cup_{i=1}^{k}c_{i})$ is maximized. A generalized version of Max Cover, where the aim is to maximize, by electing a set of subsets, any submodular set function [^3] For $C$ a finite set, a set function $f:\mathcal{P}(C)\to\mathbb{R}$ is submodular if, for any $W_{1},W_{2}\subseteq C$, we have $f(W_{1})+f(W_{2})\geq f(W_{1}\cup W_{2})+f(W_{1}\cap W_{2})$., was already studied in 1978, by [26], who proved that a greedy algorithm is $(1-1/\operatorname{\text{e}})$-approximate. However, it was not until twenty years later that the Maximum Coverage problem itself was introduced, by [16], who note NP-completeness of the unit-weight decision problem formulation of Max Cover, by reduction from Set Cover, also describe the greedy algorithm, and prove the approximation ratio of $1-1/\operatorname{\text{e}}$. In the same year, [10] gave yet another proof of the approximation ratio of the greedy algorithm, but more importantly, using a reduction from approximating Max 3SAT-5, Feige proved that for any $\epsilon>0$, Max Cover cannot be approximated in polynomial time within a ratio of $1-1/\operatorname{\text{e}}+\,\epsilon$, unless P=NP; the simple greedy algorithm is thus optimal.
[…]
In the literature on multi-winner voting, [32] make boundedness assumptions on the amount of sets an element appears in. The algorithm optimizes over the $\lceil\frac{2pk}{1-\beta}+k\rceil$ largest sets, outputs a $\beta$-approximate committee in time $\text{poly}(n,m)\cdot\binom{2pk/(1-\beta)+k}{k}$, under the assumption of a maximum number $p$ of sets any element may appear in. Under the assumption that each element appears in at least $p$ sets, the standard greedy algorithm attains an improved approximation ratio of $1-\text{e}^{-\max\left(\frac{pk}{m},1\right)}$.
Finally, a slight modification of the problem was studied by [12], who present a combinatorial $(1-1/\operatorname{\text{e}})$-approximation algorithm for Max Cover over a matroid constraint, using a non-oblivious local search that considers improvements of size 1 only. Non-oblivious refers to the algorithm using an auxiliary objective function to evaluate solutions. The paper also shows that, over a matroid constraint, oblivious local search, guided only by the problem’s objective function, achieves an approximation ratio of only $\frac{k-1}{2k-\ell-1}$, when improvements of size $\ell$ are considered. A greedy approach strands at ratio $1/2$.
Incomplete or inaccurate information
Single-winner voting with incomplete information has received some attention over the years in the computational social choice literature [5]. More recently, [15] studied proportionality in approval-based committee elections under incomplete information. They present a version of local search PAV that queries voters and, using $\mathcal{O}(mk^{6}\log k\log m)$ queries, finds a solution that satisfies Extended Justified Representation [1] and Optimal Average Satisfaction (based on the notion of average satisfaction, introduced by [33]) with high probability. Their result extends to an $\alpha$-approximate version of both axioms, in which case the query complexity decreases by a factor $k^{3}$. The same study also presents a lower bound of $\Omega(m^{11})$ for the query complexity of non-adaptive algorithms, that can not adapt their querying strategy to readily obtained information. In the field of fair division of indivisible items, [19] study the number of queries required to find a round-robin allocation. When voters’ responses are corrupted with a probability $p\in(0,1/2)$, the query complexity is lower bounded by $\Omega(nm\log(1/\delta)+m\log(m/\delta))$ and upper bounded by $\mathcal{O}(nm\log(m/\delta))$, in both of their two query models.
Our contribution
We study diversity in approval-based committee elections when information elicited from voters is incomplete or inaccurate. We measure diversity of a solution with its unit weight Max Cover, or equivalently, approval-based Chamberlin-Courant, score. For a large part, our work combines the two lines of research outlined above, and is based in particular on [12], [15] and [19]. We make four main contributions. First, we adapt the local search algorithm for Max Cover over a matroid constraint studied by [12] to an incomplete information setting, and prove an upper bound on the number of queries required to attain a $(1-1/\operatorname{\text{e}})\gamma$-approximate solution with high probability, where $\gamma$ is a parameter in $(0,1)$ (Theorem 4.10). For the original Max Cover problem (i.e. not over a matroid constraint) we study a greedy algorithm and also prove an upper bound for the number of queries required for execution (Theorem 4.4). Second, we prove, for any $\epsilon>0$ and $\gamma\in(\frac{1}{2},1)$, a lower bound for the query complexity of any non-adaptive algorithm that is $(1-1/\operatorname{\text{e}})\gamma$-approximate with probability more than $\epsilon$ (Theorem B.1). Our proof follows the proof strategy of [15, Theorem 3.1]. Third, for a setting with inaccurate information, we adapt the proof of Theorem 9 of [19], and establish a lower bound of $\Omega(nm\log(1/\delta))$ for the query complexity of finding a $(1-1/\operatorname{\text{e}})$-approximate solution with probability $1-\delta$ (Theorem 5.2). We attain an upper bound of $\mathcal{O}(nm\log(nm/\delta)$ (Theorem 5.1). Fourth, using real data from Polis, we provide a number of observations suggesting that our algorithms perform considerably better than what our worst-case analysis suggests. The observation appears to hold even in situations in which votes are both incomplete and inaccurate. For an overview of the theoretical results, see Figure 1.
Figure 1: Overview of the theoretical results. For specifics on e.g. range of parameters, or exceptions, see the theorem statements.
3 Preliminaries
An instance of an approval-based committee election problem consists of a set $V$ of $n$ voters, a set $C$ of $m$ candidates, and a natural number $k$. The goal is to elect a committee $W\subseteq C$ of size $k\leq m$, based on the opinions of the voters in $V$. In the context of deliberation summarization, we often even have $k\ll m$. An instance also contains approval information: every voter $i\in V$ approves of a subset of the candidates. Usually, this approval information is known upfront and expressed in the form of an approval set $A_{i}$ for each voter $i\in V$. The sequence of sets $A=(A_{1},\ldots,A_{n})$ is called the approval profile, and we call this setting the perfect information setting. This paper studies incomplete and inaccurate information settings; the respective models are detailed in the corresponding section.
To measure diversity of an elected committee, we use the approval-based Chamberlin-Courant score, which is equivalent to the unit weight Max Cover score:
Definition 3.1 ([7]).
On any approval-based committee election instance $(V,C,A,k)$, the approval-based Chamberlin-Courant (CC) score of a committee $W\subseteq C,\,|W|=k$ is
$\text{CC}(W)=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}_{\{A_{i}\cap W\neq\emptyset\}}(i).$ (1)
In the approval-based Chamberlin-Courant decision problem, the input is an instance $(V,C,A,k)$ and an integer $x$, and the question is whether it is possible to achieve score $\frac{x}{n}$ using $k$ candidates. In the approval-based Chamberlin-Courant optimization problem, the input is an instance $(V,C,A,k)$ and the task is to maximize the score using $k$ candidates.
[…]
Since we study only approval-based committee elections, and the approval-based Chamberlin-Courant problem is equivalent to the unit-weight Maximum Coverage problem, we use both terms interchangeably. Similarly, we refer interchangeably to voters/elements, to candidates/sets, and committees/covers. Whenever we speak about the Chamberlin-Courant problem, the approval-based version is meant, and whenever the Maximum Coverage problem is mentioned, the unit-weight version is meant.
4 Incomplete information
In the incomplete information setting, voters are only able to express their votes (approve/disapprove) on a small portion of the candidates, so we don’t have access to all approval information beforehand. To model this, we assume there exists an underlying true distribution $d$ of approval preferences, and we now write $A:V\to\mathcal{P}(C)$ for a random variable, distributed according to $d$, that maps a voter $i$ to the set $A(i)$ of candidates that voter $i$ approves of. To obtain approval information, an algorithm must draw samples from $d$. We assume that voters are able to respond to at most $t$ approve/disapprove questions at one go, so drawing voter $i$, we obtain, for some set $Q\subseteq C$ for which $|Q|=t$, the set $A(i)\cap Q$. We call the voter-candidate pair $(i,j)$ a query and $Q$ a query set. This means query set $Q$ allows us to perform $t$ queries per queried voter. We are interested in finding a solution to Max Cover in this incomplete information setting. When voters are able to express their vote on all candidates, $nm$ query responses suffice to gather all information, and we will surely find a $(1-1/\operatorname{\text{e}})$-approximate solution, e.g. with a greedy approach. When voters are not able to express their vote on all candidates, the algorithm’s performance might drop. In this section, we present theoretical results on the relation between the approximation ratio and the number of queries required to obtain that ratio. We study two settings. Firstly, the original problem, where we may elect any $k$-sized committee we want, and, secondly, the problem over a matroid constraint, where the elected committee has to be part of a matroid that is defined with the problem instance. For the proof of a lower bound for the query complexity of non-adaptive algorithms in the original problem, we refer to the appendix.
4.1 Upper bound for the unconstrained setting
The standard greedy algorithm
We prepare the ground by considering the perfect information setting. Algorithm 1 is the greedy algorithm for this setting. We write $\Delta(W,c^{\prime})$ for the increase in Chamberlin-Courant score obtained by adding $c^{\prime}$ to committee $W$. Here, but also in future algorithms, we abuse notation by using $\in$ for lists.
Input: Numbers $n,m,k\in\operatorname{\mathbb{N}}$ with $k\leq m$, set $V$ of $n$ voters, set $C$ of $m$ candidates.
Output: Committee $W\subseteq C$ of size $k$.
Let $W=[\ ]$ be an empty list;
for $i=1,\ldots,k$, do
  1. 1.
    For all $c^{\prime}\notin W$, determine $\Delta(W,c^{\prime})$.
  2. 2.
    Append $W$ with $\operatorname*{argmax}_{c^{\prime}\notin W}\,\Delta(W,c^{\prime})$.
      end for
ALGORITHM 1 Greedy
Lemma 4.1.
Algorithm 1 is $(1-\frac{1}{\operatorname{\text{e}}})$-approximate for the Max Cover, and runs in time $\mathcal{O}(kmn)$.
This is a well known result, see e.g. [16]. We provide our own proof of Lemma 4.1 in the appendix. Writing $\operatorname{\textsc{gr}}(k)$ for the Max Cover score of the greedy algorithm after termination, we will assume that $\sum_{i\in V}\mathds{1}(|A(i)|\geq 1)\geq k$; there exist at least $k$ voters that approve a candidate. We call such instances, for which $\sum_{i\in V}\mathds{1}(|A(i)|\geq 1)\geq k$ is true, non-trivial instances. After running the algorithm, thereby gathering the query responses, we can verify efficiently whether an instance is non-trivial.
Proposition 4.2.
On any non-trivial instance, $\operatorname{\textsc{gr}}(k)\geq k$.
For the incomplete information setting, that we discuss in detail in the next section, we estimate approval information of the total population, based on the information of only few voters. This can lead to errors. Write $\operatorname{\textsc{gr}}\text{-}\epsilon$ for an adaptation of Algorithm 2 that, in each iteration, might choose any candidate that yields the largest increase in score up to $-\epsilon$. We can compute the approximation ratio in a similar fashion as before in Lemma 4.1.
Lemma 4.3.
For any non-trivial instance, greedy-$\epsilon$ is $\left(\left(1-\frac{1}{\operatorname{\text{e}}}\right)\cdot\left(\frac{1}{1+\epsilon(1-\frac{1}{\operatorname{\text{e}}})}\right)\right)$-approximate.
The greedy query algorithm
We now turn to the incomplete information setting. Recall that here we assume there is an underlying unknown distribution $d$ of voter preferences and we write $A:V\to\mathcal{P}(C)$ for the random variable that maps a voter $i$ to its approval set $A(i)$, and is distributed according to $d$. In the query model, we repeatedly sample sets (of size $\ell\leq n$) of voters independently from $d$ and ask these voters for their votes on some query set $Q\subseteq C$ (of size $t$). Querying any set $Q\subseteq C$, on all voters, the responses allow us to compute, for any $S\subseteq Q$,
$p_{S}:=\mathbb{P}(A\cap S\neq\emptyset)=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}(A(i)\cap S\neq\emptyset);$ (2)
the Chamberlin-Courant score of $S$. We can now write $\Delta(W,c^{\prime}):=p_{W\cup\{c^{\prime}\}}-p_{W}$, and we can compute this with complete information on a query containing $W\cup\{c^{\prime}\}$. Sampling $\ell$ voters from $d$ independently and presenting them with query set $Q\subseteq C$, we can compute, for every subset $S\subseteq Q$,
$\hat{p}_{S}:=\frac{1}{\ell}\sum_{i=1}^{\ell}\mathds{1}(\{A(i)\cap S\neq\emptyset\}),$ (3)
that estimates $p_{S}$. We write $\hat{\Delta}(W,c^{\prime})=\hat{p}_{W\cup\{c^{\prime}\}}-\hat{p}_{W}$ for an estimate of $\Delta(W,c^{\prime})$ based on values $\hat{p}$. Using this querying approach, we obtain Algorithm 2. Theorem 4.4 gives an upper bound on the amount of query sets required to run Algorithm 2.
Input: Numbers $n,m,k\in\operatorname{\mathbb{N}}$ with $k\leq m$, set $V$ of $n$ voters, set $C\subseteq\mathcal{P}(V)$ of $m$ candidates, query size $t$, with $m\geq t>k$, and $\gamma\in(0,1),\ \delta\in(0,1]$.
Output: Committee $W\subseteq C$ of size $k$.
Set $\epsilon=\frac{(1-\gamma)e}{\gamma(e-1)}$ and $\ell=\lceil\frac{2}{\epsilon^{2}}(\log(\frac{2mk}{\delta}))\rceil$;
Let $W=[\ ]$ be an empty list;
for $i=1,\ldots,k$, do
  1. 1.
    Construct a set of query sets $\mathcal{Q}=\{Q_{i}\}_{i}$ with $Q_{i}\subseteq C$ and $|Q_{i}|=t$ for all $i$, and such that $W\subseteq\bigcap\mathcal{Q}$ and $\bigcup\mathcal{Q}=C$, and present each query set to $\ell$ voters.
  2. 2.
    For all $c^{\prime}\notin W$, determine $\hat{\Delta}(W,c^{\prime})$ as an estimate of $\Delta(W,c^{\prime})$ using $\ell$ voters and query set $Q$ containing $W\cup c^{\prime}$.
  3. 3.
    Append $W$ with $\operatorname*{argmax}_{c^{\prime}\notin W}\,\hat{\Delta}(W,c^{\prime})$.
      end for
ALGORITHM 2 Greedy-queries
Theorem 4.4.
For any non-trivial instance, any $\delta>0,\,\gamma\in(0,1)$, and $k<t\leq m$, with probability at least $1-\delta$, Algorithm 2 is $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover using
$k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil 0.8\left(\frac{\gamma}{1-\gamma}\right)^{2}\log\left(\frac{2mk}{\delta}\right)\right\rceil\in\mathcal{O}\left(\left(\frac{\gamma}{1-\gamma}\right)^{2}km\log\left(\frac{km}{\delta}\right)\right)$
queries.
Example 4.5.
For $m=1000,\,\gamma=0.85,\,\delta=0.05,\,k=8$ and $t=20$, the upper bound of Theorem 4.4 gives 432 920 queries.
4.2 Upper bound for a setting with matroid constraints
Matroids in multi-winner voting
We can imagine situations where we want to place restrictions on the set of possible committees to be elected. For instance, perhaps we want to include ‘external’ diversity requirements, such as ‘the committee must contain at least $x$ candidates of category $Y$’. In practice, $x$ would be a natural number and $Y$ would be a category like gender, race, etc. In online deliberation settings, these attributes would extend to users. Or perhaps we want to place a restriction on how many comments can be selected that are submitted by one and the same person. For such purposes, we make a selection of which committees are “allowed” to be elected, and collect them in a set $\mathcal{M}$. This is fundamentally different from what we have so far meant with diversity, which was just defined in terms of the approval profile. It is not obvious that any algorithm adapts well to this restriction. Fortunately, [12] prove that, when $\mathcal{M}$ is a matroid, we can mend this problem.
Definition 4.6 (matroid).
A matroid $\mathcal{M}$ over a set $C$ is a selection of subsets (called independent sets) that satisfies
  1. 1.
    $\emptyset\in\mathcal{M}$,
  2. 2.
    if $A\in\mathcal{M}$ and $B\subset A$ then $B\in\mathcal{M}$, and
  3. 3.
    for all $A,B\in\mathcal{M}$ with $|A|>|B|$, there exists $x\in A\setminus B$ such that $B\cup\{x\}\in\mathcal{M}$.
This last property guarantees that all maximal independent sets have the same cardinality. We call such sets bases, and their common cardinality is called the rank of the matroid.
Example 4.7 (uniform matroid).
Consider =$\mathcal{M}=\{W\subset C:|W|\leq k\}$. This is the uniform matroid of rank $k$: 1) $|\emptyset|=0\leq k$, 2) for any set $W\in\mathcal{M}$ with $|W|=j\leq k$, for any subset $W^{\prime}\subseteq W$, $|W^{\prime}|\leq|W|\leq k$, and 3) for any two feasible sets $W,\ W^{\prime}\in\mathcal{M}$ with $|W^{\prime}|<|W|$, we will have, for any $c\in W\setminus W^{\prime}$, that $|W^{\prime}\cup\{c\}|\leq k$.
Thus, taking for $\mathcal{M}$ the uniform matroid of rank $k$, we retrieve the original problem. This shows the problem with matroid constraints is a generalization of the original problem, hence more difficult.
Example 4.8 (a matroid implementing quota).
Define a set of bases $M=\{W\subseteq C:|W|=k,\ \ |W\cap Y|=\frac{1}{2}k,\ |W\cap X|=\frac{1}{2}k\}$ and consider $\mathcal{M}=M\bigcup\ \{V:V\subseteq W,\ W\in M\}$. This is a matroid of rank $k$.
In Example 4.8, only the bases elements comply with the quota. This is not a problem, we show next.
Proposition 4.9.
For any instance $(V,C,A,k)$ and any matroid $\mathcal{M}$ of rank $k$, no cover of size $\ell<k$ will attain a unique maximal Max Cover score.
Thus, there is always basis among those covers with the highest score. When an algorithm does output a cover $L$ of size $\ell<k$, we can take any basis $K$ and add elements from $K$ to $L$ until it has rank $k$.
The standard non-oblivious local search algorithm
As done above, we start with the complete information setting. Here, a classical (or, oblivious) local search algorithm (for Max Cover) typically starts with an arbitrary (tentative) solution and, in each iteration, swaps one or multiple of the tentatively elected sets for tentatively unelected sets in order to improve the objective function value. It thus searches in the local neighborhood of the solution. It stops when no local improvement is possible anymore, which could in principle take exponentially many steps. As described in [12], this running time problem can be circumvented. We can use approximate local search, which allows us to bound the running time at a small cost in the approximation ratio, see also [27]. In practice, this means we stop the algorithm sooner, namely when we encounter, in some iteration, an improvement smaller than some parameter $\beta>0$. Then, using a partial enumeration technique, we can eliminate this small cost again, see [8]. This is also the approach that [12] take.
[12] prove that, over a matroid constraint, the greedy algorithm would attain only approximation ratio $\frac{1}{2}$. Also, the approximation ratio of oblivious local search (Algorithm 1 in their paper)is $\frac{k-1}{2k-\ell-1}$ when $\ell$ sets are exchanged in each iteration. For $k=2$ and $\ell=1$, this also equals $\frac{1}{2}$, which is quite a lot smaller than $1-1/\operatorname{\text{e}}$. Thus, they switch to a non-oblivious local search algorithm that uses an auxiliary objective function $f$ for the iterative procedure. This auxiliary objective function adds temporary weight to elements covered multiple times. This makes a path of improvements more “robust”, and thereby prevents getting stuck in bad local optima. We denote by $\alpha_{j}$ the temporary weight associated to an element covered $j$ times, and write $h_{i}(W)$ for the amount of times that $i$ is covered by $W$. Then, the auxiliary objective function value of committee $W$ is
$f(W)=\frac{1}{n}\sum_{i=1}^{n}\alpha_{h_{i}(W)}.$ (4)
Note that, with $\alpha_{0}=0$ and $\alpha_{j}=1$ for all $j>0$, we retain the original oblivious objective function, but if we set $\alpha_{j}>\alpha_{1}$ for $j>1$, we add additional weight to elements covered multiple times. As in [12], we define
$\alpha_{0}=0,\ \ \ \ \ \ \ \ \alpha_{1}=1-\frac{1}{\operatorname{\text{e}}},\ \ \ \ \ \ \ \ \alpha_{j+1}=(j+1)\alpha_{j}-j\alpha_{j-1}-\frac{1}{\operatorname{\text{e}}}.$ (5)
The sequence $(\alpha_{n})_{n\in\operatorname{\mathbb{N}}_{0}}$ thus gives zero points to a voter that is not covered by the committee at all, $1-\frac{1}{\operatorname{\text{e}}}\approx 0.632$ points to a voter that is covered by a single committee member, $2-\frac{3}{\operatorname{\text{e}}}\approx 0.896$ points to a voter that is covered twice, and so on. [12] prove this choice for the sequence $(\alpha_{n})_{n\in\operatorname{\mathbb{N}}_{0}}$ is optimal and that, for any $\gamma>0$, using this auxiliary objective function, the non-oblivious local search algorithm, with parameter $\beta$ (decreasing in $\gamma$), is $(1-1/\operatorname{\text{e}}-\gamma)$-approximate.
We have adjusted the non-oblivious local search algorithm of [12] to suit our setting better. We explain the changes below (in Section 4.2), as they are induced by the querying strategy. For now it is only important to know that the changes do not affect the approximation guarantees of the algorithm. See our adapted version, Algorithm 3, below, where, for $c\in W$ and $c^{\prime}\notin W$, we now write $\Delta(W,c^{\prime}):=f(W\cup\{c^{\prime}\})-f(W)$ and $\Delta(W,c^{\prime},c):=f((W\cup\{c^{\prime}\})\setminus\{c\})-f(W)$.
Input: Numbers $n,m,k\in\operatorname{\mathbb{N}}$ with $k\leq m$, set $V$ of $n$ voters, set $C$ of $m$ candidates, $\beta>0$.
Output: Committee $W$, consisting of $k$ candidates.
Choose $W\subset C$ such that $|W|=k$, and $c\in W$ and $c^{\prime}\notin W$ so that $(W\cup\{c^{\prime}\})\setminus\{c\}\in\mathcal{M}$;
repeat
  1. 1.
    $W=(W\cup\{c^{\prime}\})\setminus\{c\}$.
  2. 2.
    Let $\mathcal{E}$ be the set of all valid exchanges for $W$ according to $\mathcal{M}$.
  3. 3.
    Pick $(c^{\prime},c)\in\operatorname*{argmax}_{(x,y)\in\mathcal{E}}\Delta(W,x,y)$.
      until $\Delta(W,c^{\prime},c)\leq\beta$ and $f(W)>1/n$;
ALGORITHM 3 Local Search-$\beta$
[12, Corollary 6] prove that the local search algorithm with parameter $\beta=\frac{\gamma}{C_{0}k\log k(1-1/\operatorname{\text{e}}-\gamma)}$, is $(1-1/\operatorname{\text{e}}-\gamma)$-approximate for any $\gamma>0$ and some universal constant $C_{0}$ for which (by their Lemma 4) we know that $\alpha_{i}\leq C_{0}\log i$ for all $i\leq k$. Simplifying notation, we write $\beta=C_{1}\frac{\gamma}{k\log k}$ for any $\gamma>0$ and some universal constant $C_{1}=\frac{1}{C_{0}(1-1/\operatorname{\text{e}}-\gamma)}\leq\frac{\log i}{\alpha_{i}(1-1/\operatorname{\text{e}}-\gamma)}$ for any $i\leq k$.
The local search query algorithm
We turn to the incomplete information setting. Like before, we write $A:V\to\mathcal{P}(C)$ for the random variable, with unknown distribution $d$, that maps a voter $i$ to approval set $A(i)$. After querying any set $Q\subseteq C$ on all voters, we can compute, for any $S\subseteq Q$,
$p_{S}:=\mathbb{P}(A\cap Q=S)=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}(A(i)\cap Q=S),$ (6)
for the probability that a voter approves, within the set $Q$, of exactly the set $S$. Note that this is different than how $p_{S}$ was defined in Section 4.1, but both are meant to denote the relevant information that we attain from a query set $Q$. With this information, we can compute $f(Q)$. Since this time, we have, for $c^{\prime}\notin W$ and $c\in W$, $\Delta(W,c^{\prime}):=f(W\cup\{c^{\prime}\})-f(W)$ and $\Delta(W,c^{\prime},c):=f((W\cup\{c^{\prime}\})\setminus\{c\})-f(W)$, we can compute these values with complete information on a query set containing $W\cup\{c^{\prime}\}$. Indeed,
$\begin{split}f(W)&=\frac{1}{n}\sum_{i=1}^{n}\alpha_{h_{i}(S)}\\ &=\frac{1}{n}\sum_{i=1}^{n}\sum_{S\subseteq W}\mathds{1}(A(i)\cap W=S)\cdot\alpha_{|S|}\\ &=\sum_{S\subseteq W}\frac{1}{n}\sum_{i=1}^{n}\mathds{1}(A(i)\cap W=S)\cdot\alpha_{|S|}\\ &=\sum_{S\subseteq W}p_{S}\cdot\alpha_{|S|}.\end{split}$ (7)
Sampling $\ell\leq n$ voters from $d$ independently and presenting them with query set $Q\subseteq C$, we can compute, for every subset $S\subseteq Q$,
$\hat{p}_{S}:=\frac{1}{\ell}\sum_{i=1}^{\ell}\mathds{1}(A(i)\cap Q=S),$ (8)
that estimates $p_{S}$. Consequently, we write $\hat{\Delta}(W,c^{\prime})=\hat{f}(W\cup\{c^{\prime}\})-\hat{f}(W)=\sum_{S\subseteq W\cup\{c^{\prime}\}}\hat{p}_{S}\cdot\alpha_{|S|}-\sum_{S\subseteq W}\hat{p}_{S}\cdot\alpha_{|S|}$ for an estimate of $\Delta(W,c^{\prime})$ based on the values $\hat{p}$) and $\hat{\Delta}(W,c^{\prime},c)$ as an estimate of $\Delta(W,c^{\prime},c)$ based on values $\hat{p}$ analogously. We present our querying algorithm in Algorithm 4.
Input: Numbers $n,m,k\in\operatorname{\mathbb{N}}$ with $k\leq m$, set $V$ of $n$ voters, set $C$ of $m$ candidates, $\beta>0$, query size $t$, with $m\geq t>k$, constants $\delta\in(0,1]$ and $\xi\geq 1$.
Output: Committee $W$, consisting of $k$ candidates.
$\epsilon=\frac{\xi-1}{2\xi}\cdot\beta$, $\ell=\left\lceil\frac{(2-2/\operatorname{\text{e}})^{2}}{2\epsilon^{2}}\log\left(\frac{2\cdot(m-k)k\cdot\xi\alpha_{k}}{\delta\cdot\beta}\right)\right\rceil$;
Choose $W\subset C$ such that $|W|=k$, and $c\in W$ and $c^{\prime}\notin W$ so that $(W\cup\{c^{\prime}\})\setminus\{c\}\in\mathcal{M}$;
repeat
  1. 1.
    $W=(W\cup\{c^{\prime}\})\setminus\{c\}$;
  2. 2.
    Let $\mathcal{E}$ be the set of all valid exchanges for $W$ according to $\mathcal{M}$.
  3. 3.
    Construct a set of query sets $\mathcal{Q}=\{Q_{i}\}_{i}$ with $Q_{i}\subseteq C$ and $|Q_{i}|=t$ for all $i$, and such that $W\subseteq\bigcap\mathcal{Q}$ and $\bigcup\mathcal{Q}=C$, and present each query set to $\ell$ voters.
  4. 4.
    For all $c\in W,\,c^{\prime}\notin W$, determine $\hat{\Delta}(W,c^{\prime},c)$ as an estimate of $\Delta(W,c^{\prime},c)$ using $\ell$ voters and query set $Q$ containing $W\cup\{c^{\prime}\}$.
  5. 5.
    Pick $(c^{\prime},c)\in\operatorname*{argmax}_{(x,y)\in\mathcal{E}}\hat{\Delta}(W,x,y)$.
      until $\hat{\Delta}(W,c^{\prime},c)<\beta-\epsilon$ and $\hat{f}(W)>\alpha_{1}/n$;
ALGORITHM 4 Local Search-$\beta$-queries
As previously mentioned, in Algorithm 3 (hence also Algorithm 4) we implemented several changes with respect to Algorithm 2 of [12]. We explain them here. There are three changes. Firstly, we have left out the greedy initialization. Although it speeds up the convergence time-wise, it would cost many extra querying rounds. It is not possible to query only for the initialization and use the gathered information for the local search phase, since candidates not elected in the initialization, or only elected in its last round, will not have query responses from all voters whereas that is necessary if the candidate would become part of the tentative solution at some point during the local search. However, it might be possible to save the information gathered in the initialization in a smart way. For now, we leave out this phase. For similar reasons, we omit the partial enumeration step that would yield the clean $1-1/\operatorname{\text{e}}$ ratio. Secondly, we have adapted the stopping condition from $\frac{f((W\cup c^{\prime})\setminus c)}{f(W)}\leq 1+\beta n$ (the factor $n$ is added because we use relative scores) to $f((W\cup c^{\prime})\setminus c)-f(W)\leq\beta$. Before we can make any kind of statement about the performance of our algorithm, we need guarantees on the accuracy of our estimations. In order to use Hoeffding’s inequality to bound the probability that our estimates are far from the estimated value, we need our samples to be bounded as well. However, since $\alpha_{0}=0$, the expression $\frac{f((W\cup c^{\prime})\setminus c)}{f(W)}$ is unbounded. Suppose now, that $f((W\cup c^{\prime})\setminus c)-f(W)\leq\beta$, then
$f((W\cup c^{\prime})\setminus c)-f(W)\leq\beta\implies\frac{f((W\cup c^{\prime})\setminus c)}{f(W)}\leq 1+\frac{\beta}{f(W)},$ (9)
which is at most $1+\beta n$ iff $f(W)\geq 1/n$. This might not be true immediately, but will be true at some point as long as $\sum_{i\in V}\mathds{1}(|A(i)|\geq 1)\geq 1$, so also when the instance is non-trivial. Thus, we also add the requirement that $\hat{f}(W)\geq\alpha_{1}/n$, and then our adapted stopping condition is a stronger requirement for any non-trivial instance. This makes our third adaptation. Theorem 4.10 gives the amount of queries required for Algorithm 4. As expected, this amount is higher than stated in Theorem 4.4 for the original problem.
Theorem 4.10.
For any $\delta>0$, $\gamma\in(0,1)$, $m\geq t>k$, and some universal constant $C_{2}$, with probability at least $1-\delta$, Algorithm 4, with $\beta=C_{2}\frac{1-\gamma}{\gamma k\log k}$, is $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover when using
$\xi\alpha_{k}\cdot\left(\frac{\gamma k\log k}{C_{2}(1-\gamma)}\right)^{3}\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil\frac{(2-2/\operatorname{\text{e}})^{2}}{2}\left(\frac{2\xi}{\xi-1}\right)^{2}\log\left(\frac{2\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\gamma k\log k}{C_{2}(1-\gamma)\delta}\right)\right\rceil$
queries. Asymptotically in $k,m,\gamma$ and $\delta$, this is in
$\mathcal{O}\left(\left(\frac{\gamma k\log k}{1-\gamma}\right)^{3}m\log\left(\frac{m\gamma k^{2}\log k}{\delta(1-\gamma)}\right)\right).$
Example 4.11.
The number of queries increases in $\xi$ when $\xi\geq 3$, and has a negative asymptote at $\xi=1$, so regardless of the values of the other parameters, we attain a minimum for some $\xi\in[1,3]$. Fixing $\xi=3$, for $m=1000,\,\delta=0.05,\,k=8$ and $t=20$ the upper bound of Theorem 4.10 gives $9.52180\cdot 10^{11}$ queries; quite a lot.
5 Inaccurate information
In the inaccurate information setting, voters do hand in complete ballots, but the voters are inaccurate in reporting their true approvals: each time we query a voter $i$ about a candidate $j$ (so, posing query $(i,j)$), the response is corrupted with probability $p$. We model this by adding, to the original complete information approval profile $A$, a random variable $X\sim\text{Bernoulli}(p)$, with $p\in(0,\frac{1}{2})$. Posing query $(i,j)$, we also draw from $\text{Bernoulli}(p)$, and the result determines whether we see a truthful response (draw 0) or not (draw 1). Draws from $X$ are independent between voters, between candidates, and between consecutive draws for the same query. If $p=0$, we would have accurate information, and with $nm$ queries we would have all, accurate, information. We would then be sure to find a $(1-1/\operatorname{\text{e}})$-approximate solution (e.g. using the greedy algorithm). Now $p>0$, so without adapting the algorithms, their performance will drop. We here present theoretical results on the amount of queries required to still obtain a $(1-1/\operatorname{\text{e}})$-approximate solution with high probability.
When answers may be corrupted, we can simply pose every question multiple times to compensate for uncertainty. That is the strategy we use to obtain the upper bound. We can adapt Algorithm 1 (attaining Algorithm 5 , see Appendix) or Algorithm 3; it doesn’t matter, since the new total amount of queries only depends, besides on $p$ and $\delta$, on the original amount of queries, which would be $nm$ for both algorithms. This also means we immediately have an upper bound for the problem over a matroid constraint.
Theorem 5.1.
In the query model with inaccurate query responses and parameter $p\in(0,\frac{1}{2})$, for any $\delta>0$, $n,m\in\mathbb{N}$, we attain a $(1-1/\operatorname{\text{e}})$-approximate solution to Max Cover with probability at least $1-\delta$ using $nm\cdot U$ queries, where $U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$. This gives a query complexity in $\mathcal{O}(nm\log(nm/\delta))$.
We prove the lower bound using an information theoretic lower bound from multi-armed bandit theory. The proof disregards the possibility of matroid constraints, but since the original problem is a special case of the problem over a matroid constraint, the lower bound holds for both settings.
Theorem 5.2.
In the query model with inaccurate query responses, for any $\delta\in(0,1),\,n,m\in\mathbb{N}$, any algorithm that outputs an ($1-1/\operatorname{\text{e}}$)-approximate solution to Max Cover with probability at least $1-\delta$, makes $\Omega(nm\log(1/\delta))$ queries in expectation.
6 Experiments
Our experimental section is motivated by three questions. First, although our theoretical analysis (Theorems 4.4 and 4.10) provides good asymptotic bounds on the query complexity of both Greedy and Local Search-$\beta$ (Algorithms 2 and 4), large constant factors are a big influence for relatively small instances. For both theorems, the number of voters $\ell$ that needs to be sampled for each query set in order to arrive at the stated approximation ratio with high probability, is very large. This is not only impractical (each voter would need to reply more than $n$ times) but also makes the computations extremely slow. However, as the theoretical bounds are worst-case, we might be able to attain good results even on smaller instances. Second, Theorem 5.1 provides a good asymptotic bound on the number of queries required to compensate inaccurate voting behavior, requiring a number $U\in\mathcal{O}(nm/\delta)$ of repetitions for each query. We are interested in the experimental performance of our algorithms when we ignore this advice (that is, take $U=1$). Third and finally, we are interested in the performance of our algorithms when we combine incomplete and inaccurate information. We did not study theoretically in this paper, but a naive approach for an upper bound would be to combine both separate bounds (by multiplication). What can we say about this experimentally?
We investigate these questions by running both our querying algorithms (Algorithms 2 and 4) and their exact versions (Algorithms 1 and 3) on 18 open-use datasets provided by Polis  [^5] https://github.com/compdemocracy/openData, using $p=0$ and $p=\frac{1}{10}$. We used a committee size $k=8$ in all experiments (changing $k$ resulted in very similar results).
(a) Mean and standard deviation of the Max Cover score of Greedy (Algorithm 1), Local Search-$\beta$ (Algorithm 3) and Approval Voting on the 18 Polis datasets at each iteration of the algorithms for $k=8$. Since the score of Local Search-$\beta$ depends on the randomly sampled initial committee, we ran 50 trials of that algorithm.
(b) Mean and standard deviation of the Max Cover score of Greedy-queries (Algorithm 2) and Local Search-queries (Algorithm 4) on the 18 Polis datasets at each iteration of the algorithms for $k=8$. We ran 50 trials of both algorithms, as the performance depends on which voters are randomly sampled, as well as on the initial committee, for Local Search-queries. The score is expressed as a fraction of the score of the exact version of the algorithms on the given data set.
Figure 2: Performance of our algorithms, mean and standard deviation taken over 18 Polis datasets
We start with the full information case. In Figure 2(a) we show the Max Cover score, obtained by Greedy (Algorithm 1) and Local Search-$\beta$ (Algorithm 3), per iteration. We took $\beta$ as in Theorem 4.10 (and will write Local Search going forward). We also show the Max Cover score that Approval Voting (selecting the $k$ most popular statements) would reach. Note that Greedy and Approval Voting each have a committee of size $i$ in iteration $i$, whereas Local Search starts with a random committee of size $k$, swapping one committee member each iteration. This explains the differences in scores for low iteration numbers. After $k=8$ iterations, we see that the score is confidently above $1-1/\operatorname{\text{e}}\approx 0.63$, indicating that the approximation ratio is achieved, even when the optimal score is 1. We also see that Greedy and Local Search perform better than Approval Voting, as expected.
Taking the scores of the Greedy and Local Search from Figure 2(a) as our baseline, we want to understand how close we can get to these scores when using only few queries, and/or when information is inaccurate. Recall that $t$ denotes the number of statements contained in a single query set. Taking $t=20$, we write $M$ for the expected number of query sets of size $t$ that each voter is given. We then configure Local Search-queries (Algorithm 4), and Greedy-queries (Algorithm 2) so that $M=1,2,3,4,5$. This is far lower than what would be theoretically required. We do this for both $p=0$ and $p=\frac{1}{10}$. Figure 2(b) shows the performance of both algorithms relative to their exact versions.
We see that both algorithms with accurate responses ($p=0$) obtain a score very close ($0.9-0.95$) to the exact versions of the algorithms, even when $M=1$. For the same number of queries, the local search algorithm seems to perform slightly closer to its exact version than the greedy algorithm does. Note that the exact versions of both algorithms obtained an average Max Cover score of around $0.73$. Multiplying by $0.9$ still yields a score above the worst-case approximation ratio of $1-1/\operatorname{\text{e}}\approx 0.63$, even when the optimal solution would have a Max Cover score of 1. Example 6.1 highlights the exceptional performance witnessed in the experiments for $p=0$.
Example 6.1.
The dataset vtaiwan.uberx has $m=197$ and $n=1921$. In Figure 2(b), we see that, for $M=5$, the Max Cover score attained by Local Search-queries, on average across 18 data sets (and 20 runs per set) is approximately 0.95 of the score attained by Local Search. So, for our example, let’s take $\gamma=0.95$. Then, for $k=8$ and $t=20$, taking $\delta=0.05$, we would need $M=7.109\cdot 10^{8}$.
Taking $p=\frac{1}{10}$, but with complete information, we make a similar observation. For both algorithms, adding an inaccuracy factor $p=\frac{1}{10}$ decreases the performance with a factor $0.95-0.97$ on average. Taking into consideration the performance of the exact versions of the algorithms, we see, like before, a score confidently above the worst-case approximation ratio of $1-1/\operatorname{\text{e}}$, whereas our theoretical results required taking multiple samples for each query $(i,j)$ to compensate for the inaccuracy. Example 6.2 shows a computation of the theoretically required values.
Example 6.2.
Theorem 5.1 provided us with a constant factor to multiply the number queries with in order to guarantee a $(1-1/\operatorname{\text{e}})$-approximate solution with probability $\delta>0$. We compute $U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$, and taking again as example vtaiwan.uberx, with $m=197$ and $n=1921$. Then, for $\delta=0.05$, $U=32$. For our experiments, we used $U=1$.
Taking both $p=\frac{1}{10}$ and incomplete information, not meeting the theoretically required number repetitions for both scenarios, the performance takes a noticeable hit. These query algorithms perform at around $0.75-0.85$ of the exact versions of the algorithm, with a much bigger standard deviation. Remember, however, that we only queried a handful of statements per voter, and voters were possibly inaccurate, whereas the exact versions of the algorithms have complete access to all accurate information.
7 Conclusion
We studied approval-based committee elections with the goal of maximizing the Chamberlin-Courant score under incomplete or inaccurate information. The study of this type of multi-winner elections, and the focus on diversity, are motivated by applications in digital democracy for the support of online large-scale deliberation.
Analytically, the paper determines bounds on the query complexity for the above settings, in particular: for the incomplete information setting, we establish upper bounds for adaptations to our setting of known algorithms for the Max Cover problem (Theorems Theorem 4.10 and 4.4) and a lower bound (Theorem B.1) for any non-adaptive algorithm. For the inaccurate information setting, we establish a lower bound for optimal approximations with high probability (Theorem 5.2), as well as an upper bound using a greedy algorithm (Theorem 5.1). Empirically, we complement the worst-case results by testing the behavior of the algorithms on augmented real-world data from Polis. Even when voters respond, possibly inaccurately, to just a handful of statements, our algorithms perform well compared to the algorithms with complete and accurate access to each voters’ ballots.
A first direction for future work is to query voters adaptively, to reduce the amount of queries required. Second, we hope to combine both incomplete and inaccurate information theoretically. Third, we wish to develop a more comprehensive theory of diversity, beyond the reliance on the Chamberlin-Courant score. Fourth, we aim to strengthen our experimental section by including synthetic data. Fifth, we intend to explore the significance of the above results also to applications in active learning [31], with the aim of optimizing the querying of a group annotators in distributed data labeling tasks (see, e.g., [22]).
Acknowledgments.
This research was partially supported by Hybrid Intelligence Center, a 10-year programme funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://www.hybrid-intelligence-centre.nl/, under Grant No. (024.004.022), and by the European Union under the Horizon Europe grant Perycles (Participatory Democracy that Scales, http://perycles-project.eu/).
Appendix A Proofs of Section 4
Lemma 4.1.
Algorithm 1 is $(1-\frac{1}{\operatorname{\text{e}}})$-approximate for the maximum coverage problem, and runs in time $\mathcal{O}(kmn)$.
For the proof we use the following notation: for $i\leq k$, let $\operatorname{\textsc{gr}}_{i}$ denote the increase in Max Cover score the greedy algorithm attains in iteration $i$. Also for $i\leq k$, write $\operatorname{\textsc{gr}}(i)$ for the Max Cover score of the greedy solution after $i$ iterations, so $\operatorname{\textsc{gr}}(k)$ is the final score. We write $\operatorname{\textsc{opt}}(k)$ for the score an optimal solution, consisting of $k$ candidates, would have.
Proof.
We start with the approximation ratio. Like [16, Lemma 1], we first prove that, for $\ell\leq k$, $\operatorname{\textsc{gr}}_{\ell}\geq(\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1))/k$. Fix an optimal solution. At the start of iteration $\ell$, the difference between the greedy score and the optimal score is $\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1)$. Thus, there are at least this many elements covered in the (fixed) optimal solution, that have not been covered yet by greedy up to that point (it could be more: the set selected by greedy at the start of iteration $\ell$, needn’t be a subset of the optimal solution). These $\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1)$ elements are covered by the $k$ sets in the optimal solution so, by the pigeon hole principle, there exists a set in the optimal solution, that is not currently selected by greedy, that will cover at least $(\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1))/k$ of these elements. Then, in iteration $\ell$, the greedy algorithm picks the set that covers the most additional elements, so these must be at least $(\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1))/k$ many. We use this relation to obtain an expression for $\operatorname{\textsc{gr}}(k)$ in terms of $\operatorname{\textsc{opt}}(k)$. For all $k$,
$\operatorname{\textsc{gr}}(k)$ $=\operatorname{\textsc{gr}}(k-1)+\operatorname{\textsc{gr}}_{k}$
$\geq\operatorname{\textsc{gr}}(k-1)+\frac{\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(k-1)}{k}$
$=\frac{\operatorname{\textsc{opt}}(k)}{k}+\operatorname{\textsc{gr}}(k-1)\left(1-\frac{1}{k}\right)$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}+\operatorname{\textsc{gr}}(k-2)\left(1-\frac{1}{k}\right)\right)\left(1-\frac{1}{k}\right)$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}+\operatorname{\textsc{gr}}(k-3)\left(1-\frac{1}{k}\right)\right)\left(1-\frac{1}{k}\right)\right)\left(1-\frac{1}{k}\right)$
$\ldots$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}+\frac{\operatorname{\textsc{opt}}(k)}{k}\left(1-\frac{1}{k}\right)+\frac{\operatorname{\textsc{opt}}(k)}{k}\left(1-\frac{1}{k}\right)^{2}+\ldots+\frac{\operatorname{\textsc{opt}}(k)}{k}\left(1-\frac{1}{k}\right)^{k-2}+\operatorname{\textsc{gr}}(1)\left(1-\frac{1}{k}\right)^{k-1}$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}\left(\sum_{i=0}^{k-2}\left(1-\frac{1}{k}\right)^{i}+\left(1-\frac{1}{k}\right)^{k-1}\right)$
$=\frac{\operatorname{\textsc{opt}}(k)}{k}\sum_{i=0}^{k-1}\left(1-\frac{1}{k}\right)^{i}$
$=\frac{\operatorname{\textsc{opt}}(k)}{k}\left(k\left(1-\left(1-\frac{1}{k}\right)^{k}\right)\right)$
$=\operatorname{\textsc{opt}}(k)\left(1-\left(1-\frac{1}{k}\right)^{k}\right).$
Because $\left(1-\frac{1}{k}\right)^{k}$ is an increasing sequence with $\lim_{k\to\infty}\left(1-\frac{1}{k}\right)^{k}=\frac{1}{\operatorname{\text{e}}}$, we can bound the last expression from below by $\operatorname{\textsc{opt}}(k)\left(1-\frac{1}{\operatorname{\text{e}}}\right)$, which proves the approximation ratio.
Regarding the time complexity, the algorithm chooses $k$ candidates sequentially. For each of these $k$ steps, we have to go over all $m$ candidates to check which adds most to the current score. Per candidate, we have to check for each of the $n$ voters whether it is already represented (we can keep track of the represented voters in one set). This gives a time complexity in $\mathcal{O}(kmn)$. ∎
Remark A.1.
The convergence goes rather quickly; for $k=10$, we already have $1-(1-\frac{1}{k})^{k}=1-\frac{9}{10}^{10}\approx 0.651$, whereas $1-1/\operatorname{\text{e}}\approx 0.632$. However, for small committee sizes the actual worst case approximation ratio can be substantially smaller than the general ratio of $1-1/\operatorname{\text{e}}$: when $k=3$, the ratio is actually $1-\frac{2}{3}^{3}\approx 0.704$.
Proposition 4.2.
On any non-trivial instance, $\operatorname{\textsc{gr}}(k)\geq k$.
Proof.
We proceed by contraposition. Assume $\operatorname{\textsc{gr}}(k)<k$, which is equivalent to $\operatorname{\textsc{gr}}(k)\leq k-1$, so (by the pigeonhole principle) there exists $i\leq k$ such that $\operatorname{\textsc{gr}}_{i}=0$. Because $\operatorname{\textsc{gr}}_{i}\leq\operatorname{\textsc{gr}}_{i-1}$ for all $i\leq k$, we must have (at least) $\operatorname{\textsc{gr}}_{k}=0$; no candidate with a yet uncovered voter is left in iteration $k$, since otherwise the greedy algorithm would pick that over an empty candidate and then $\operatorname{\textsc{gr}}_{k}>0$. At this point, $\operatorname{\textsc{gr}}(k-1)\leq k-1$ and no uncovered voters are left, so $\sum_{i\in V}\mathds{1}(|A(i)|\geq 1)\leq k-1$. ∎
Lemma 4.3.
For any non-trivial instance, greedy-$\epsilon$ is $\left(\left(1-\frac{1}{\operatorname{\text{e}}}\right)\cdot\left(\frac{1}{1+\epsilon(1-\frac{1}{\operatorname{\text{e}}})}\right)\right)$-approximate.
For $i\leq k$, define $\operatorname{\textsc{gr}}\text{-}\epsilon_{i}$ and $\operatorname{\textsc{gr}}\text{-}\epsilon(i)$ similarly to before.
Proof.
We write out $\operatorname{\textsc{gr}}\text{-}\epsilon(k)$ recursively like before, but now subtracting $\epsilon$ in each iteration. For any $k$, we have
$\operatorname{\textsc{gr}}\text{-}\epsilon(k)$ $=\operatorname{\textsc{gr}}\text{-}\epsilon(k-1)+\operatorname{\textsc{gr}}\text{-}\epsilon_{k}$
$\geq\operatorname{\textsc{gr}}\text{-}\epsilon(k-1)+\frac{\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}\text{-}\epsilon(k-1)}{k}-\epsilon$
$=\frac{\operatorname{\textsc{opt}}(k)}{k}+\operatorname{\textsc{gr}}\text{-}\epsilon(k-1)\left(1-\frac{1}{k}\right)-\epsilon$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}+\operatorname{\textsc{gr}}\text{-}\epsilon(k-2)\left(1-\frac{1}{k}\right)-\epsilon\right)\left(1-\frac{1}{k}\right)-\epsilon$
$=\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)+\operatorname{\textsc{gr}}\text{-}\epsilon(k-2)\left(1-\frac{1}{k}\right)^{2}$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)^{2}+\operatorname{\textsc{gr}}\text{-}\epsilon(k-3)\left(1-\frac{1}{k}\right)^{3}$$\ldots$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)^{2}$
$\ \ \ +\ldots+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)^{k-2}+\operatorname{\textsc{gr}}\text{-}\epsilon(1)\left(1-\frac{1}{k}\right)^{k-1}$
$\geq\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\sum_{i=0}^{k-1}\left(1-\frac{1}{k}\right)^{i}$
$=\operatorname{\textsc{opt}}(k)\left(1-\left(1-\frac{1}{k}\right)^{k}\right)-\epsilon k\left(1-\left(1-\frac{1}{k}\right)^{k}\right),$
which we can bound from below by $(\operatorname{\textsc{opt}}(k)-\epsilon k)\left(1-\frac{1}{\operatorname{\text{e}}}\right)$ again by taking a limit $k\to\infty$. This shows $\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)}{\operatorname{\textsc{opt}}(k)}\geq\left(1-\frac{\epsilon k}{\operatorname{\textsc{opt}}(k)}\right)\left(1-\frac{1}{\operatorname{\text{e}}}\right)=\left(1-\frac{1}{\operatorname{\text{e}}}\right)+\frac{1}{\operatorname{\textsc{opt}}(k)}\left(\frac{\epsilon k}{\operatorname{\text{e}}}-\epsilon k\right)\iff\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)+\epsilon k\left(1-\frac{1}{\operatorname{\text{e}}}\right)}{\operatorname{\textsc{opt}}(k)}\geq 1-\frac{1}{\operatorname{\text{e}}}.$
Using $\operatorname{\textsc{gr}}(k)\geq k$, we can reformulate $1-\frac{1}{\operatorname{\text{e}}}\leq\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)+\epsilon k\left(1-\frac{1}{\operatorname{\text{e}}}\right)}{\operatorname{\textsc{opt}}(k)}\leq\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)+\operatorname{\textsc{gr}}\text{-}\epsilon(k)\cdot\epsilon\left(1-\frac{1}{\operatorname{\text{e}}}\right)}{\operatorname{\textsc{opt}}(k)}=\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)\cdot\left(1+\epsilon\left(1-\frac{1}{\operatorname{\text{e}}}\right)\right)}{\operatorname{\textsc{opt}}(k)}$. ∎
Theorem 4.4.
For any non-trivial instance, any $\delta>0,\,\gamma\in(0,1)$, and $k<t\leq m$, with probability at least $1-\delta$, Algorithm 2 is $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover using
$k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil 0.8\left(\frac{\gamma}{1-\gamma}\right)^{2}\log\left(\frac{2mk}{\delta}\right)\right\rceil\in\mathcal{O}\left(\left(\frac{\gamma}{1-\gamma}\right)^{2}km\log\left(\frac{km}{\delta}\right)\right)$
queries.
Proof.
We will compute the necessary values of $\epsilon$ and $\ell$ so that the statement is true, and we will see that they are indeed as stated in Algorithm 2. From Proposition 4.2 and Lemma 4.3, we know that if all our estimates are within $\frac{1}{2}\epsilon$ of the value that they estimate, then $\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)}{\operatorname{\textsc{opt}}(k)}\geq\left(1-\frac{1}{\operatorname{\text{e}}}\right)\left(\frac{1}{1+\epsilon(1-\frac{1}{\operatorname{\text{e}}})}\right)$. For $\gamma\in(0,1)$, we take $\epsilon=\frac{(1-\gamma)e}{\gamma(e-1)}$ so that $\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)}{\operatorname{\textsc{opt}}(k)}\geq\gamma\cdot\left(1-\frac{1}{\operatorname{\text{e}}}\right)$.
We want to find $\ell$ so that with probability $1-\delta$, all our estimations $\hat{\Delta}$ are within $\frac{1}{2}\epsilon$ of the true values $\Delta$. For this, we use Hoeffding’s inequality to bound the probability that a single estimate is more than $\frac{1}{2}\epsilon$ away from its corresponding true value. We assume our samples are independent, and we know that all samples are in $\{0,1\}$, that their sample mean is $\hat{\Delta}$ and their expected value is $\Delta$, so using Hoeffding’s inequality we obtain for any estimate $\hat{\Delta}$ of $\Delta$:
$\mathbb{P}\left(|\hat{\Delta}-\Delta|\geq\frac{1}{2}\epsilon\right)\leq 2\exp\left(-\frac{1}{2}\ell\epsilon^{2}\right).$ (10)
There are $k$ iterations and at most $m$ estimates per iteration so no more than $k\cdot m$ estimations in total. Thus, by independence, we have
$\mathbb{P}\left(\forall i:|\hat{\Delta}_{i}-\Delta_{i}|<\frac{1}{2}\epsilon\right)\geq 1-2\exp\left(-\frac{1}{2}\ell\epsilon^{2}\right)\cdot m\cdot k.$ (11)
We want to find $\ell$ so that this value equals at least $1-\delta$, and so we proceed to write
$1-2\exp\left(-\frac{1}{2}\ell\epsilon^{2}\right)\cdot m\cdot k\geq 1-\delta\iff\exp\left(-\frac{1}{2}\ell\epsilon^{2}\right)\leq\frac{\delta}{2mk}\iff\ell\geq\frac{2}{\epsilon^{2}}\log\left(\frac{2mk}{\delta}\right).$ (12)
This means, for any $0<\delta\leq 1$, any $\ell$ that is at least this large, will, with probability $1-\delta$, produce (querying $\ell$ voters) estimates that are all at most $\frac{1}{2}\epsilon$ apart from the value that they estimate. Then we can invoke Lemma 4.3 to obtain the desired ratio. We take the smallest possible value: $\ell=\left\lceil\frac{2}{\epsilon^{2}}\log\left(\frac{2mk}{\delta}\right)\right\rceil$ and filling in for $\epsilon$ this results in
$\ell=\left\lceil\frac{2}{\left(\frac{(1-\gamma)e}{\gamma(e-1)}\right)^{2}}\log\left(\frac{2mk}{\delta}\right)\right\rceil=\left\lceil\frac{2\gamma^{2}(e-1)^{2}}{(1-\gamma)^{2}\operatorname{\text{e}}^{2}}\log\left(\frac{2mk}{\delta}\right)\right\rceil\leq\left\lceil 0.8\left(\frac{\gamma}{1-\gamma}\right)^{2}\log\left(\frac{2mk}{\delta}\right)\right\rceil\ ,$ (13)
giving
$k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\ell\leq k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil 0.8\left(\frac{\gamma}{1-\gamma}\right)^{2}\log\left(\frac{2mk}{\delta}\right)\right\rceil\in\mathcal{O}\left(\left(\frac{\gamma}{1-\gamma}\right)^{2}km\log\left(\frac{km}{\delta}\right)\right)$ (14)
queries. ∎
Proposition 4.9.
For any instance $(V,C,A,k)$ and any matroid $\mathcal{M}$ of rank $k$, no cover of size $\ell<k$ will attain a unique maximal Max Cover score.
Proof.
Suppose there does exist a cover $L\in\mathcal{M}$ of size $\ell<k$ that has a higher Max Cover score than any size-$k$ cover $M\in\mathcal{M}$. This means there exists no superset of $L$ in $\mathcal{M}$, in particular not any of size $k$, because the Max Cover score is monotone in the amount of sets that the cover contains. But by requirement 3) in Definition 4.6, using any basis element $K\in\mathcal{M}$, we can make a path of sets $L,\ L\cup k_{1},\ L\cup k_{1}\cup k_{2},\ \ldots,L\cup k_{1}\cup\ldots\cup k_{k-\ell}=:L^{\prime}$ from $L$ to attain a basis element $L^{\prime}\in\mathcal{M}$ that contains $L$. ∎
Theorem 4.10.
For any $\delta>0$, $\gamma\in(0,1)$, $m\geq t>k$, and some universal constant $C_{2}$, with probability at least $1-\delta$, Algorithm 4, with $\beta=C_{2}\frac{1-\gamma}{\gamma k\log k}$, is $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover when using
$\xi\alpha_{k}\cdot\left(\frac{\gamma k\log k}{C_{2}(1-\gamma)}\right)^{3}\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil\frac{(2-2/\operatorname{\text{e}})^{2}}{2}\left(\frac{2\xi}{\xi-1}\right)^{2}\log\left(\frac{2\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\gamma k\log k}{C_{2}(1-\gamma)\delta}\right)\right\rceil$
queries. Asymptotically in $k,m,\gamma$ and $\delta$, this is in
$\mathcal{O}\left(\left(\frac{\gamma k\log k}{1-\gamma}\right)^{3}m\log\left(\frac{m\gamma k^{2}\log k}{\delta(1-\gamma)}\right)\right).$
Proof.
We will compute the necessary values of $\epsilon$ and $\ell$ so that the statement is true, and we will see that they are indeed the same as stated in Algorithm 4. In the original approximation ratio (of Algorithm 3), we substitute $\gamma$ for $(1-1/\operatorname{\text{e}})(1-\gamma)$, so that the ratio becomes $(1-1/\operatorname{\text{e}})\cdot\gamma$ and we adapt the constant factor accordingly.
With this value of $\beta$, the minimum step size of Local Search-$\beta$ (Algorithm 3) is $C_{2}\frac{1-\gamma}{\gamma k\log k}$, and every committee attains score at least 00 and at most $\alpha_{k}$, so Algorithm 3 runs at most $\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}$ iterations. We anticipate that querying will increase the amount of required iterations, because of the uncertainty inducing estimations. Depending on how many iterations we allow, values of $\epsilon$ (the estimation margin) and $\ell$ (the amount of voters queried per round) will also differ. Thus, to make this dependency explicit, we add in a factor $\xi\geq 1$ and say the algorithm with queries has a total of at most $\xi\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}$ iterations. We then compute the maximal value of $\epsilon$ this corresponds to.
Until the last run of the algorithm, we have $\hat{\Delta}(W,c^{\prime},c)>\beta-\epsilon$, and when this is no longer true, the swap of $c$ for $c^{\prime}$ is not made. We take $\epsilon$ so that all improvements ${\Delta}(W,c^{\prime},c)$ are at least size $\hat{\Delta}(W,c^{\prime},c)-\epsilon>\beta-2\epsilon$ and solve
$\beta-2\epsilon=C_{2}\frac{1-\gamma}{\gamma k\log k}-2\epsilon=C_{2}\frac{1-\gamma}{\xi\cdot\gamma k\log k},$ (15)
for $\epsilon$ to obtain $\epsilon=\frac{\xi-1}{2\xi}\cdot C_{2}\frac{1-\gamma}{\gamma k\log k}$, which is the maximum margin of estimation we can allow in order to have our original number of iterations multiplied by at most $\xi$.
We want to have $\ell$ such that with probability $1-\delta$, all our estimations $\hat{\Delta}$ are within $\epsilon$ of the true values $\Delta$. To this end, we use Hoeffding’s inequality to bound the probability that a single estimation is more than $\epsilon$ away from its corresponding $\Delta$-value. For any committee, any voter will attain some score in $\{\alpha_{0},\ldots,\alpha_{k}\}$ and whenever a candidate $c\in W$ is replaced for another candidate $c^{\prime}\notin W$, the voter can approve one less candidate than before, the same amount of candidates as before, or one more candidate than before. A score $\alpha_{i}$ before a swap, can thus become any of scores $\{\alpha_{i-1},\alpha_{i},\alpha_{i+1}\}$ after. In Lemma 3, [12] prove that for all $i<k$, $\alpha_{i+1}>\alpha_{i}$ and $\alpha_{i+2}-\alpha_{i+1}\leq\alpha_{i+1}-\alpha_{i}$. Thus, $\{\alpha_{i}\}_{i}$ is an increasing sequence, but the increase declines monotonically. We may thus take $\alpha_{1}-\alpha_{0}=1-\frac{1}{\operatorname{\text{e}}}$ for an upper bound on the difference between consecutive values in $\{\alpha_{i}\}_{i}$. Then we see that all our samples of $\Delta(W,c)$ and $\Delta(W,c^{\prime},c)$ take values in the interval $[-1+\frac{1}{\operatorname{\text{e}}},1-\frac{1}{\operatorname{\text{e}}}]$. The sample mean is $\hat{\Delta}$ and the expected value is $\Delta$, and we assume the samples are independent. Hoeffding’s inequality then yields, for any of the estimates $\hat{\Delta}$ of $\Delta$,
$\mathbb{P}(|\hat{\Delta}-\Delta|\geq\epsilon)\leq 2\exp\left(-\frac{2\epsilon^{2}\ell}{(2-2/\operatorname{\text{e}})^{2}}\right).$ (16)
In each iteration we estimate $\Delta(W,c^{\prime},c)$ based on all combinations $(c^{\prime},c)$, so these are $(m-k)\cdot k$ estimations. This gives $(m-k)\cdot k\cdot\xi\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}$ estimations in total over all iterations. The probability that for all $i$, the estimation $\hat{\Delta}_{i}$ of $\Delta_{i}$ is within $\epsilon$ of the true value, then becomes
$\mathbb{P}(\forall i:|\hat{\Delta}_{i}-\Delta_{i}|<\epsilon)\geq 1-2\exp\left(-\frac{2\epsilon^{2}\ell}{(2-2/\operatorname{\text{e}})^{2}}\right)\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}.$ (17)
We can now choose $\ell$ such that this it at least $1-\delta$:
$\begin{split}1-2&\exp\left(-\frac{2\epsilon^{2}\ell}{(2-2/\operatorname{\text{e}})^{2}}\right)\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}\geq 1-\delta\\ \iff&2\exp\left(-\frac{2\epsilon^{2}\ell}{(2-2/\operatorname{\text{e}})^{2}}\right)\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}\leq\delta\\ \iff&\exp\left(-\frac{2\epsilon^{2}\ell}{(2-2/\operatorname{\text{e}})^{2}}\right)\leq\frac{1}{2}\delta\cdot\frac{1}{(m-k)k}\cdot\frac{1}{\xi\alpha_{k}}\cdot\frac{C_{2}(1-\gamma)}{\gamma k\log k}\\ \iff&-\ell\leq\frac{(2-2/\operatorname{\text{e}})^{2}}{2\epsilon^{2}}\log\left(\frac{1}{2}\delta\cdot\frac{1}{(m-k)k}\cdot\frac{1}{\xi\alpha_{k}}\cdot\frac{C_{2}(1-\gamma)}{\gamma k\log k}\right)\\ \iff&\ell\geq\frac{(2-2/\operatorname{\text{e}})^{2}}{2\epsilon^{2}}\log\left(\frac{2\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\gamma k\log k}{C_{2}(1-\gamma)\delta}\right).\end{split}$ (18)
This means, for any $0<\delta\leq 1$, any $\ell$ that is at least this size, will, with probability $1-\delta$, produce (querying $\ell$ voters) estimates that are all at most $\epsilon$ apart from the value that they estimate. We take the smallest possible value of $\ell$:
$\ell=\left\lceil\frac{(2-2/\operatorname{\text{e}})^{2}}{2\epsilon^{2}}\log\left(\frac{2\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\gamma k\log k}{C_{2}(1-\gamma)\delta}\right)\right\rceil.$
The algorithm terminates when $\hat{\Delta}(W,c^{\prime},c)<\frac{C_{2}(1-\gamma)}{\gamma k\log k}-\epsilon$, which implies $\Delta(W,c^{\prime},c)<\frac{C_{2}(1-\gamma)}{\gamma k\log k}$. This is the original stopping condition of the algorithm without queries, so the original performance guarantees also hold here; the approximation ratio is $(1-1/\operatorname{\text{e}})\gamma$. Last we study the amount of queries. In total there are
$\xi\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil\frac{(2-2/\operatorname{\text{e}})^{2}}{2\epsilon^{2}}\log\left(\frac{2\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\gamma k\log k}{C_{2}(1-\gamma)\delta}\right)\right\rceil$ (19)
$=\xi\alpha_{k}\cdot\left(\frac{\gamma k\log k}{C_{2}(1-\gamma)}\right)^{3}\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil\frac{(2-2/\operatorname{\text{e}})^{2}}{2}\left(\frac{2\xi}{\xi-1}\right)^{2}\log\left(\frac{2\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\gamma k\log k}{C_{2}(1-\gamma)\delta}\right)\right\rceil$
queries. Asymptotically in $k\ll m,\,\gamma$ and $\delta$, this is in $\mathcal{O}\left(\left(\frac{\gamma k\log k}{1-\gamma}\right)^{3}m\log\left(\frac{m\gamma k^{2}\log k}{\delta(1-\gamma)}\right)\right)$. ∎
Appendix B Lower bound for the query complexity of non-adaptive algorithms in the incomplete information setting
For the incomplete information setting, we have seen two algorithms. Both algorithms make use of information acquired during the run of the algorithm. Indeed, when selecting the query sets, we need these to contain the currently elected committee. We next present a lower bound for the query complexity of non-adaptive algorithms, which can not adapt their querying strategy to readily obtained information. The proof concerns the original setting, without matroid constraints, but the result extends to the problem over a matroid constraint, as it is a more general problem. The proof strategy follows [15, Theorem 3.1].
Theorem B.1.
For any $k\geq 2$, $\epsilon>0$, and $\gamma\in(\frac{1}{2},1)$, any non-adaptive committee selection algorithm that makes fewer than $\Omega(\frac{m^{2}\epsilon}{1-\gamma})$ queries, attains a $(1-1/\operatorname{\text{e}})\gamma$-approximate max cover score with probability at most $\epsilon$.
Proof.
[…]
For any party $P_{i}$ consider any set $S\subset C_{i}$ of size $\ell$ that is covered by a query set $Q$ of size $t$. As soon as $\ell>t$ (so when $m$ is large enough), we can split $S$ into two parts: a part of size $2\leq j\leq t$ that is actually covered by $Q$, and a part of size $\ell-j$ that is outside of $Q$. The number of sets $S\subset C_{i}$ of size $\ell$ that a single query can cover, then becomes
$\sum_{j=2}^{t}\binom{t}{j}\binom{\lfloor(m-r)/p\rfloor-t}{\ell-j}\leq 2^{t}\binom{m/p}{\ell-2}\leq 2^{t}\left(\frac{m}{p}\right)^{\ell-2}$ (20)
if $\ell-2\leq\frac{m}{2p}$, because $\binom{n}{k}$ grows in $n$, and grows in $k$ until $k=n/2$. Note that we require $m$ to be large enough for this. With our $Um^{2}$ queries, at most
$Um^{2}\cdot 2^{t}\left(\frac{m}{p}\right)^{\ell-2}=Um^{\ell}2^{t}p^{2-\ell}$ (21)
sets of size $\ell$ within $C_{i}$ can be covered in total. For $m$ large enough, there are
$\binom{\lfloor(m-r)/p\rfloor}{\ell}\geq\frac{(\lfloor(m-r)/p\rfloor-\ell)^{\ell}}{\ell!}\geq\frac{1}{\ell!}\left(\frac{m}{2p}\right)^{\ell}$ (22)
[…]
$\frac{\ell!(2p)^{\ell}Um^{\ell}2^{t}p^{2-\ell}}{m^{\ell}}=\ell!2^{\ell+t}Up^{2}.$ (23)
[…]
(alg selects at least $\mathbb{P}(\operatorname{\textsc{alg}}\text{ selects at least }q\text{ designated candidates})=\sum_{j=q}^{p}\mathbb{P}(\operatorname{\textsc{alg}}\text{ selects }j\text{ designated candidates})$ (24)
$\leq\sum_{j=q}^{p}\sum_{h=0}^{j}\mathbb{P}\left(\{\operatorname{\textsc{alg}}\text{ selects }j-h\text{ designated candidates by chance}\}\bigcap\,\{\operatorname{\textsc{alg}}\text{ covers }h\text{ of the sets }S_{i}\}\right),$
where the inequality is due to the fact that covering a set $S_{i}$ is necessary to distinguish its designated candidate, but might not be sufficient. For $h=0$, we have bounded the probability above as $p/\ell$. Each of the other at most $(p-q)p$ probabilities is bounded from above by the probability of covering one set $S_{i}$, which is at most $\ell!2^{\ell+t}Up^{2}$. If we fix $U=\frac{\delta}{\ell!2^{t+\ell}p^{3}(p-q)}$ this becomes at most $\delta$. The total probability is then at most $\frac{p}{\ell}+\delta$. Recall that we want this probability to be at most $\epsilon$. We can achieve this, for instance, by fixing both terms to be $\epsilon/2$. Note that we choose $\delta$ (as it determines the query complexity), so it suffices to show that $\frac{p}{\ell}\leq\frac{\epsilon}{2}$. Recall that $p=2k-1$, so that this condition in met when $l\geq\frac{k-1}{\epsilon}$.
Recall that the query complexity was $Um^{2}$ and that $U=\frac{\epsilon/2}{\ell!2^{t+\ell}p^{3}(p-q)}$, where $p=(k-1)/2$. We want to denote the scaling of the complexity in terms of $\gamma$ and $\epsilon$. Recall that $p=(k-1)/2$ and $q=\left\lceil\left(1-\frac{1}{\operatorname{\text{e}}}\right)\gamma+\left(\left(1-\frac{1}{\operatorname{\text{e}}}\right)2\gamma-1\right)p-1\right\rceil$. We then get
$\frac{1}{p-q}\leq\frac{2}{1+k-2(1+((1-\frac{1}{\operatorname{\text{e}}})\gamma-\frac{1}{2})k)}=\frac{2}{2k(1-(1-\frac{1}{\operatorname{\text{e}}})\gamma)-1}\leq\frac{1}{k(1-\gamma)},$ (25)
where the last step required $22k\gamma/2\geq 1/$ 2, which is true in case $k\geq 2,\gamma>1/2$. It is not hard to see that $\frac{1}{\ell!2^{t+\ell}p^{3}}\leq 1$, so that the final query complexity becomes $\frac{m^{2}\epsilon}{1-\gamma}$. ∎
Theorem B.1 motivates the study of adaptive algorithms. Ideally, we would have more precise non-asymptotic results, especially because the upper bound of Algorithm 4 contains large constants. Or otherwise, it might be possible to extend this result for a larger exponent of $m$ (as done in [15]), by adapting the instance. Moreover, obviously, a lower bound for adaptive setting is desirable. Both are left for future work.
Appendix C Additions to Section 5
Input: Numbers $n,m,k\in\operatorname{\mathbb{N}}$ with $k\leq m$, set $V$ of $n$ voters, set $C$ of $m$ candidates, parameters $p\in(0,\frac{1}{2}),\,\delta>0$.
Output: Committee $W\subseteq C$ of size $k$.
$U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$;
For all $i\in V,\,j\in C$, perform query $(i,j)$ $U$ times and take the majority winner as the ‘true’ query response;
Let $W=[\ ]$ be an empty list;
for $i=1,\ldots,k$, do
  1. 1.
    For all $c^{\prime}\notin W$, determine $\Delta(W,c^{\prime})$.
  2. 2.
    Append $W$ with $\operatorname*{argmax}_{c^{\prime}\notin W}\,\Delta(W,c^{\prime})$.
      end for
ALGORITHM 5 Greedy-inaccurate
Theorem 5.1.
In the query model with inaccurate query responses and parameter $p\in(0,\frac{1}{2})$, for any $\delta>0$, $n,m\in\mathbb{N}$, we attain a $(1-1/\operatorname{\text{e}})$-approximate solution to Max Cover with probability at least $1-\delta$ using $nm\cdot U$ queries, where $U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$. This gives a query complexity in $\mathcal{O}(nm\log(nm/\delta))$.
Proof.
We adapt Algorithm 1: for all pairs $(i,j)$ for $i\in V$ and $j\in C$, we perform query $(i,j)$ $U$ times and take the majority winner of the results to be the ‘true’ value, breaking ties arbitrarily. There are $nm$ pairs $(i,j)$ for $i\in V$ and $j\in C$, so we only need to show that for any $\delta>0$, taking $U$ as stated, we obtain an $(1-1/\operatorname{\text{e}})$-approximate solution with probability $1-\delta$.
Take $\Omega=\{0,1\}^{U}$, $\mathcal{F}=\mathcal{P}(\Omega)$ and $\mathbb{P}:\mathcal{F}\to[0,1]$ that satisfies ($\mathbb{P}(\{\mathbf{v}\})=(1-p)^{y}p^{U-y}$ when $\mathbf{v}$ contains $y$ 1s and $U-y$ 0s. Accordingly, write $\mathbf{v}_{i,j}$ for a vector that is the outcome of $U$ throws of $(i,j)$, where entry 0 means the outcome is correct (success) and 1 means the outcome is corrupted (failure). Then take $X:\Omega\to\operatorname{\mathbb{N}}:\mathbf{v}_{i,j}\mapsto\sum_{l=1}^{U}v_{i,j,l}$ for the random variable that maps an outcome vector to the number of successes and write $\bar{X}:=U-X$ for the number of failures.
A Chernoff bound gives $\mathbb{P}(\bar{X}\geq\frac{1}{2}U)\leq\exp(-UD(\frac{1}{2}||p))$ [14] and we can reformulate to
$\begin{split}\exp\left\{-UD\left(\frac{1}{2}||p\right)\right\}&=\exp\left\{-U\left(\frac{1}{2}\log\left(\frac{1/2}{1-p}\right)+\frac{1}{2}\log\left(\frac{1/2}{p}\right)\right)\right\}\\ &=\exp\left\{-U\frac{1}{2}\log\left(\frac{1/2}{1-p}\right)\right\}\exp\left\{-U\frac{1}{2}\log\left(\frac{1/2}{p}\right)\right\}\\ &=\left(\frac{1}{4(1-p)p}\right)^{-\frac{1}{2}U}.\end{split}$ (26)
This gives an upper bound on the probability that we have an incorrect majority for a single $(i,j)$, where each throw is correct with probability $1-p$. We choose $U$ so that this stays smaller than $\frac{\delta}{nm}$, and for that we need
$-\frac{1}{2}U\geq\log_{\frac{1}{4p(1-p)}}\left(\frac{\delta}{nm}\right)=\frac{\log\frac{\delta}{nm}}{\log\frac{1}{4p(1-p)}}=-\frac{\log\frac{nm}{\delta}}{\log\frac{1}{4p(1-p)}},$ (27)
so we can take $U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$. Then the probability of any of the $nm$ estimates being incorrect, is bounded above by $\delta$. ∎
Theorem 5.2.
In the query model with inaccurate query responses, for any $\delta\in(0,1),\,n,m\in\mathbb{N}$, any algorithm that outputs an ($1-1/\operatorname{\text{e}}$)-approximate solution to Max Cover with probability at least $1-\delta$, makes $\Omega(nm\log(1/\delta))$ queries in expectation.
To establish the theorem, we adapt the proof of Theorem 9 in [19], using a result from multi-armed bandit theory [20] to lower bound the query complexity. The Kullback-Leibler divergence between two probability distributions is defined as $D(P||Q)=\sum_{x}P(x)\log\frac{P(x)}{Q(x)}$ for discrete probability functions $P$ and $Q$, as long as, for all $x,\,Q(x)=0\implies P(x)=0$ (otherwise we define it as +$\infty$+ ∞). This reduces to $d(a,b)=a\log\frac{a}{b}+(1-a)\log\frac{1-a}{1-b}$ for two Bernoulli distributions with parameters $a,b\in(0,1)$.
Proof.
[…]
Lemma C.1.
Let $\nu$ and $\nu^{\prime}$ be any two bandit instances defined on the same set of arms $\mathcal{A}$, with corresponding observation distributions $\{P_{a}\}_{a\in\mathcal{A}}$ and $\{P^{\prime}_{a}\}_{a\in\mathcal{A}}$. Let $\tau$ be the total number of queries made when the algorithm terminates, and let $\mathcal{E}$ be any probabilistic event that can be deduced from the resulting history $(a_{1},y_{1},\ldots,a_{\tau},y_{\tau})$, possibly with additional randomness independent of that history. Then, for $T_{a}$ the number of times action $a$ is queried up to termination index $\tau$, we have
$\sum_{a\in\mathcal{A}}\operatorname{\mathbb{E}}^{\nu}[T_{a}]D(P_{a}||P^{\prime}_{a})\geq d(\mathbb{P}^{\nu}(\mathcal{E}),\mathbb{P}^{\nu^{\prime}}(\mathcal{E})).$ (28)
For any $k\geq 2$, consider an instance $\nu$ with $m$ candidates and $n$ voters. We partition the voters into $m-w$ disjoint “parties" $P_{1},\ldots,P_{m-w}$, where $w\geq 0$ and $w$ is bounded from above by some fraction of $m$. Voters in party $P_{i}$ approve precisely $c_{i}$. Candidates $c_{1},\ldots,c_{m-w-k}$ (together $C_{1}$) all have one approving voter. These $m-w-k$ voters (together $V_{1}$) are alone in their party. Candidates $c_{m-w-k+1},\ldots,c_{m-w}$ (together $C_{2}$) all have $u>\frac{1}{1-\frac{1}{\operatorname{\text{e}}}\cdot k}$ approving voters (together $V_{2}$), so that the only $(1-\frac{1}{\operatorname{\text{e}}})$-approximate solution consists of the $k$ candidates in $C_{2}$:
$(k-1)u+1<\left(1-\frac{1}{\operatorname{\text{e}}}\right)ku\iff\left(\frac{1}{\operatorname{\text{e}}}k-1\right)u<-1\iff u>\frac{1}{1-\frac{k}{\operatorname{\text{e}}}}.$ (29)
The remaining $w$ candidates (together $C_{3}$) have no approving voters. Finally we have $V_{4}$, a set of voters that approve of no candidate (homeless voters). For any $m,n$, we can find constants $D,E\in(0,1)$ such that $|C_{1}|\geq D\cdot m$ (which means we need $w\leq m-Dm-k$) and $(1-E)n-ku\geq|C_{1}|$ so that $|V_{4}|\geq E\cdot n$, so let’s pick any such constant $D$ and $E$. Now consider the instance $\nu^{\prime}$ that is obtained by taking, $x:=\lceil k(\frac{u}{1-1/\operatorname{\text{e}}}-1)\rceil$ homeless voters and distributing them (as) evenly (as possible) over $k$ candidates $c_{1},\ldots,c_{k}\in C_{1}$. Other than that, all preferences remain unchanged. With this value of $x$, we have $x+k>\frac{ku}{1-1/\operatorname{\text{e}}}$, which means that $\nu^{\prime}$ has a unique $(1-1/\operatorname{\text{e}})$-approximate solution in $c_{1},\ldots,c_{k}\subseteq C_{1}$, whereas the unique $(1-1/\operatorname{\text{e}})$-approximate solution of $\nu$ was the set $C_{2}$. Since both instances have only one optimal solution, and these differ, we have $L_{\nu}\cap L_{\nu^{\prime}}=\emptyset$.
(a) Instance $\nu$
(b) Instance $\nu^{\prime}$
Figure 3: When $k=2$, both instances have only one $(1-1/\operatorname{\text{e}})$-approximate solution and these differ.
[…]
So now we can rewrite the inequality in Lemma C.1 as
$\left(\sum_{i=1}^{x}\sum_{j=1}^{k}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]\right)\cdot d(1-p,p)\geq d(\delta,1-\delta),$ (30)
or equivalently, $\left(\sum_{i=1}^{x}\sum_{j=1}^{k}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]\right)\geq\frac{d(\delta,1-\delta)}{d(1-p,p)}$. We can make such an adapted instance $\nu^{\prime}$ for at least $\lfloor\frac{En}{x}\rfloor$ disjoint sets of $x$ voters in $V_{4}$ (we gather these sets in a set we call $V_{4,x}$) and at least $\lfloor\frac{Dm}{k}\rfloor$ disjoint sets of $k$ candidates in $C_{1}$ (we gather these sets in a set we call $C_{1,k}$):
$\sum_{C_{1,k}}\sum_{V_{4,x}}\left(\sum_{i=1}^{x}\sum_{j=1}^{k}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]\right)\geq\left\lfloor\frac{En}{x}\right\rfloor\left\lfloor\frac{Dm}{k}\right\rfloor\frac{d(\delta,1-\delta)}{d(1-p,p)}\geq\left\lfloor\frac{E}{x}\right\rfloor\left\lfloor\frac{D}{k}\right\rfloor\cdot mn\cdot\frac{d(\delta,1-\delta)}{d(1-p,p)}.$ (31)
We rewrite the left hand side as
$\sum_{C_{1,k}}\sum_{V_{4,x}}\left(\sum_{i=1}^{x}\sum_{j=1}^{k}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]\right)\leq\sum_{i\in V}\sum_{j\in C}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]=\operatorname{\mathbb{E}}^{\nu}\left[\sum_{i\in V}\sum_{j\in C}T_{(i,j)}\right]=\operatorname{\mathbb{E}}^{\nu}[\tau],$ (32)
so we rephrase to $\operatorname{\mathbb{E}}^{\nu}[\tau]\geq\left\lfloor\frac{E}{x}\right\rfloor\left\lfloor\frac{D}{k}\right\rfloor\cdot mn\cdot\frac{d(\delta,1-\delta)}{d(1-p,p)}$. For the right hand side, note that $p$ is a fixed constant in $(0,\frac{1}{2})$ and $d(\delta,1-\delta)\in\Omega(\log(1/\delta))$ since $\delta\leq\frac{1}{2}-c$, hence, this instance $\nu$ serves to illustrate the lower bound. ∎
Appendix D Discussion on experiments (Section 6)
We have pre-processed the Polis data before use. First, of all 20 accessible Polis data sets, we remove the london.youth.policing dataset due to a low number of participants, and the bg2025-volunteers dataset because it contains far more statements than voters. Second, as anticipated, participating voters typically vote on only a small portion of the statements, so we obtain a sparse comment-participant-matrix. Since these discussions have been completed already, we will not be able to query voters adaptively, so, for the sake of running our querying algorithms, we need to artificially complete the data. We set all unanswered voter-candidate pairs to disapprove. Third, we removed statements with more than half of voters approving, as they made attaining a high Max Cover score trivial in some datasets.